a student has money in 3 accounts that pay 5%, 7%, and 8% in annual interest. She has 3 times as much invested at 8% as she does at 5% If the total amount she has invested is $1600 and her interest for the year comes to $115, how much money does she have in each account?
Answer by mananth(16949)

a student has money in 3 accounts that pay 5%, 7%, and 8% in annual interest. She has 3 times as much invested at 8% as she does at 5% If the total amount she has invested is $1600 and her interest for the year comes to $115, how much money does she have in each account?
..
5%-----------$x
8%-----------$3x
7%-----------$(1600-4x)
..
Interest = $115
..
sum of individual interests = 115
..
0.05x+0.08*3x+0.07*(1600-4x)=115
multiply by 100
5x+8*3x+7*(1600-4x)=11500
5x+24x+11200-28x=11500
x+11200=11500
x=11500-11200
x= 300 at 5%
300*3= 900 @8%
1600-1200=400 @7%
CHECK
0.05*300+0.08*900+0.07*400= 115